### v1

# Phase Separation in Kitaev Chain

Abstract
Kitaev chain is a one-dimensional spinless fermion model that has $p$-wave superconducting (SC) states and Majorana zero modes at the edge. Usually this model is analyzed by taking only SC order parameter (OP) into account, but the situation significantly changes when OPs other than the SCOP are included. It turns out that the SC state in the latter case is prone to phase separation for moderate to strong attractiive interactions.
Kitaev chain is a one-dimensional spinless fermion system which is intensively studied as a model to realize $p$-wave superconducting (SC) states with Majorana zero modes at the edge. [1] This model is usually analyzed by taking only SC order parameter (OP) into account[2, 3, 4, 5, 6], since the original modeling started with a proximity induced superconducting state in a nanowire placed on a three dimensional superconductor.
In this study we consider a one-dimensional spinless fermion model with attractive nearest-neighbor (NN) interactions by employing a mean-field approximation with Hartree-Fock-Gorkov type decoupling. Our purpose is to consider more general situations beyond the modeling of Kitaev[1], and show that different behaviors can emerge when we include OPs other than SCOP.
We treat the model with the Hamiltonian
$\displaystyle H=$ $\displaystyle-t\sum_{j}(a^{\dagger}_{j+1}a_{j}+a^{\dagger}_{j}a_{j+1})-\mu\sum_{j}a^{\dagger}_{j}a_{j}$ (1)
$\displaystyle-V\sum_{j}a^{\dagger}_{j}a_{j}a^{\dagger}_{j+1}a_{j+1},$
where $a_{j}$ is the operator for spinless fermions and $\mu$ denotes the chemical potential. $t$ and $V$ are the transfer integral and the attractive interaction between NN sites, respectively. We decouple this Hamiltonian by taking the following OPs; $p$-wave SCOP $\Delta=-i\langle a_{j+1}a_{j}\rangle$, bond OP (BOP) $\chi=\langle a^{\dagger}_{j+1}a_{j}\rangle$, and fermion density $n=\langle a^{\dagger}_{j}a_{j}\rangle$. We impose the periodic boundary condition, and assume that $\chi$ and $\Delta$ are real. The self consistency equations can be obtained by minimizing the free energy as follows
$\displaystyle n=\frac{1}{2}-\frac{1}{2N}\sum_{k}\tanh\Big(\frac{E_{k}}{2T}\Big)\frac{\xi_{k}}{E_{k}}$ (2)
$\displaystyle\chi=-\frac{1}{2N}\sum_{k}\tanh\Big(\frac{E_{k}}{2T}\Big)\frac{\xi_{k}}{E_{k}}\cos k$ (3)
$\displaystyle\Delta=\frac{\Delta}{N}\sum_{k}\tanh\Big(\frac{E_{k}}{2T}\Big)\frac{V\sin^{2}k}{E_{k}}$ (4)
where $\xi_{k}=-2(t-V\chi)\cos k-\mu-2nV$, $\Delta_{k}=2V\Delta\sin k$, and $E_{k}=\sqrt{\xi_{k}^{2}+|\Delta_{k}|^{2}}$. NN is the total number of lattice sites.
For numerical calculations we take $t=1$ as the unit of energy, and choose $n=0.3$. (Results for other values of $n$ are qualitatively the same.) In Fig.1, the SCOP ($\Delta$) and the BOP ($\chi$) are shown for several choices of $V$ as functions of $T$. As seen the superconducting transition temperature $T_{C}$ naturally increases as $V$ becomes large.
Figure 1: (Color online) SCOP $\Delta$ and BOP $\chi$ for $t=1$ and $n=0.3$. (a) $\Delta$ for $V=1.5$, (b) $\Delta$ for $V=2.0$, (c) $\Delta$ for $V=2.5$, (d) $\chi$ for $V=1.5$, (e) $\chi$ for $V=2.0$, and (f) $\chi$ for $V=2.5$.
However, there is a hidden transion to phase separatied states as we will show by calculating the compressibility $\partial n/\partial\mu$. By taking derivatives of Eqs. (3)-(5) with respect to $\mu$, the simultaneous equations for $\partial n/\partial\mu$, $\partial\chi/\partial\mu$, and $\partial\Delta/\partial\mu$ are obtained as
$\displaystyle{\hat{A}}\left(\begin{array}[]{cc}\displaystyle\frac{\partial n}{\partial\mu}\\ \displaystyle\frac{\partial\chi}{\partial\mu}\\ \displaystyle\frac{\partial\Delta}{\partial\mu}\end{array}\right)=\left(\begin{array}[]{cc}c_{1}\\ c_{2}\\ c_{3}\end{array}\right),$ (11)
where $^{\hat{A}}$ is a $3\times 3$ matrix whose matrix elements are given as $A_{11}=1-\frac{V}{N}\sum_{k}F_{1}(k)$, $A_{12}=\frac{V}{N}\sum_{k}\cos kF_{1}(k)$, $A_{13}=\frac{2V^{2}\Delta}{N}\sum_{k}\sin^{2}kF_{2}(k)\xi_{k}$, $A_{21}=-A_{12}$, $A_{22}=1+\frac{V}{N}\sum_{k}\cos^{2}kF_{1}(k)$, $A_{23}=\frac{2V^{2}\Delta}{N}\sum_{k}\cos k\sin^{2}kF_{2}(k)\xi_{k}$, $A_{31}=-\frac{2V^{2}}{N}\sum_{k}\sin^{2}kF_{2}(k)\xi_{k}$, $A_{32}=\frac{2V^{2}}{N}\sum_{k}\cos k\sin^{2}kF_{2}(k)\xi_{k}$, and $A_{33}=\frac{4V^{3}\Delta}{N}\sum_{k}\sin^{4}kF_{2}(k)$. The constants $C_{i}$ ($i=1,2,3$) are given as $C_{1}=\frac{1}{2N}\sum_{k}F_{1}(k)$, $C_{2}=\frac{1}{2N}\sum_{k}\cos kF_{1}(k)$, and $C_{3}=\frac{V}{N}\sum_{k}\sin^{2}kF_{2}(k)\xi_{k}$, where the functions $F_{1}(k)$ and $F_{2}(k)$ are defined as $F_{1}(k)=\Delta_{k}^{2}\tanh(E_{k}/2T)/E_{k}^{3}+\xi_{k}^{2}/(2TE_{k}^{2}\cosh^{2}(E_{k}/2T))$ and $F_{2}(k)=1/(2TE_{k}^{2}\cosh^{2}(E_{k}/2T))-\tanh(E_{k}/2T)/E_{k}^{3}$, respectively. By solving Eq. (5), $\partial n/\partial\mu$, etc. can be obtained.
The $T$ dependences of the compressibility are shown in Fig.2.
Figure 2: (Color online) Temperature dependence of the compressibility $\partial n/\partial\mu$ for several choices of $V$ with $t=1$ and $n=0.3$.
For $V=1.5$, $\partial n/\partial\mu$ is positive and finite so that the spatially homogeneous state is stable. In the case of $V=2.0$ ($V=2.5$), however, $\partial n/\partial\mu$ diverges at $T\sim 0.40$ ($T\sim 0.78$), and becomes negative for $T\lesssim 0.40$ ($T\lesssim 0.78$). These results indicate that there are regions where the homogeneous states are not stable and phase separation (PS) occurs. In numerical calculations, if we take $n$ as an input parameter a spurious solution may be obtained, but the compressibility diverges (or becomes negative) in these cases. When the resultant $\mu$ is taken as an input parameter, we cannot obtain the original value of $n$ (=0.3=0.3), since multiple values of $n$ corresponds to a single value of $\mu$.
Figure 3: (Color online) Temperature dependence of the compressibility $(\partial n/\partial\mu)_{0}$ when OPs other than the SCOP are omitted. Parameters are the same as in Fig.2.
For comparison, we calculate the compressibility when only the SCOP is taken into account. This amounts to replace $\xi_{K}$ with $2t\cos k-\mu$, and we denote the compressibility in this case as $(\partial n/\partial\mu)_{0}$. In Fig.3, $T$ dependences of $(\partial n/\partial\mu)_{0}$ are shown. It is seen that $(\partial n/\partial\mu)_{0}$ is always positive and finite so that PS is absent. The temperature at which $(\partial n/\partial\mu)_{0}$ has a peak corresponds to $T_{C}$ for each $V$. Here we note that $T_{C}$ in this case is different from those in Fig.1, because the dispersion $\xi_{k}$ is diffrent.
Thie above considerations indicate that the occurence of PS is due to the OPs usually ignored. If we discard the term $2nV$ and retain the term $V\chi$, PS would not occur. This means that the former term, which is proportoinal to the particle density in the dispersion relation $\xi_{k}$, is the primary cause of PS. On the contrary, if we drop the term $V\chi$ and keep the term $2nV$, PS may occur but the value of $V$ necessary to induce PS becomes large. This is because the inclusion of $V\chi$ in $\xi_{k}$ reduce the band width, and so the density of states at the fermi level is increased, leading to the situation favorable to the phase transition. Then dropping the term $V\chi$ works to suppress PS.
Similar behavior is also observed in strongly correlated electron systems; the two-dimensional $t-t^{\prime}-J$ model treated within a slave-boson mean-field approximation[7, 8] is an example. In this case, the dispersion relation of spinons depends on the hole density and the BOP, and the contributions from their derivatives with respect to the chemical potential may lead to the divergence of the compressibility, and hence phase separation.[9]
In real systems, there exist long-range (or screened) Coulomb interactions (LRC) between electrons, and it works to suppress phase separation. It might be expected that competition between the tendency toward inhomogeneous states and the LRC would result in charge density wave states with finite wave vectors $\bf q$}, instead of ${\bf q}={\bf 0}$ in the case of PS, as a compromise between the two. This will be an intriguing future problem.
In summary, we have examined the stability of the SC states in Kitaev chain and found that it is prone to phase separarion when the attractive interaction becomes strong. Phase separation in this case is caused by the particle-density dependence of the dispersion relation, which are arising from the OPs other than SCOP but are often ignored. This issue may also be important for the stability of the superconducting states in other kinds of models.
{acknowledgment}
The author thanks M. Hayashi for useful discussions. This work was supported by JSPS KAKENHI Grant Number 23K04525.